# Enhanced Generative Recommendation via Content and Collaboration Integration

### v1

@@ Abstract.
Generative recommendation has emerged as a promising paradigm aimed at augmenting recommender systems with recent advancements in generative artificial intelligence. This task has been formulated as a sequence-to-sequence generation process, wherein the input sequence encompasses data pertaining to the user’s previously interacted items, and the output sequence denotes the generative identifier for the suggested item. However, existing generative recommendation approaches still encounter challenges in (i) effectively integrating user-item collaborative signals and item content information within a unified generative framework, and (ii) executing an efficient alignment between content information and collaborative signals.
In this paper, we introduce content-based colla-borative generation for recommender systems, denoted as ColaRec. To capture collaborative signals, the generative item identifiers are derived from a pretrained collaborative filtering model, while the user is represented through the aggregation of interacted items’ content. Subsequently, the aggregated textual description of items is fed into a language model to encapsulate content information. This integration enables ColaRec to amalgamate collaborative signals and content information within an end-to-end framework. Regarding the alignment, we propose an item indexing task to facilitate the mapping between the content-based semantic space and the interaction-based collaborative space. Additionally, a contrastive loss is introduced to ensure that items with similar collaborative GIDs possess comparable content representations, thereby enhancing alignment. To validate the efficacy of ColaRec, we conduct experiments on three benchmark datasets. Empirical results substantiate the superior performance of ColaRec.
Recommender System, Generative Recommendation, Collaborative Filtering, Generative Model, Language Model
isbn: ccs: Information systems Recommender systemsccs: Information systems Retrieval models and rankingccs: Information systems Novelty in information retrieval
@@
Recommender systems are widely deployed to discover user interests and provide personalized information services (Hu et al., 2018; Qi et al., 2020; Yuan et al., 2019). Recently, generative models, such as large language models (LLMs) (Zhao et al., 2023; Chang et al., 2023) and diffusion models (Yang et al., 2023; Cao et al., 2022), have gained prominence in advancing artificial intelligence. In such a context, generative recommendation (Wang et al., 2023b; Bao et al., 2023; Si et al., 2023; Rajput et al., 2023; Geng et al., 2022; Ji et al., 2023) has emerged as a new paradigm to enhance recommendation agents with end-to-end generation capability of these models.
Generative Recommendation. Generative recommendation assigns each item with a unique sequence of tokens as the item’s generative identifier, a.k.a., GID, then the recommendation is formulated as a sequence-to-sequence generation task, where the input is historical user-item interactions while the output sequence refers to the GID of the recommended item (Rajput et al., 2023; Si et al., 2023; Hua et al., 2023) [^1] Wang et al. (2023a) proposed another paradigm to directly generate new content, e.g., images, for recommendation. However, their methods are tailored for the generation of virtual content and cannot be used for recommendation of concrete items, such as the item recommendation in an e-commerce platform. In this paper, we target on the recommendation of concrete items..
@@
Compared with conventional itemIDs with an assigned single random token, the sequential tokens of a GID contain more explicit representation information. As shown in Figure 1, we see correlated GIDs denote correlated items, and thus help the recommender to gain better performance. Generative recommendation provides an end-to-end paradigm to effectively utilize generative models for recommendation. Tay et al. (2022) initially introduced Differentiable Search Indexing (DSI), which marks an early endeavor in employing generative models for information retrieval. DSI has inspired a series of studies on generative recommendation (Rajput et al., 2023; Si et al., 2023; Petrov and Macdonald, 2023). The promising results of these works demonstrate the potential of generative recommendation.
Limitations of Current Approaches. Existing generative recommendation approaches still have the following two limitations, leading to suboptimal performance.
Firstly, existing methods fail to effectively model collaborative signals and item content information in a unified framework. Collaborative signals refer to the knowledge contained in the user-item interactions while item content information refers to the textual description of items, as shown in Figure 2. On one hand, several methods utilize either the item title or hierarchical content embeddings of items to construct GIDs (Li et al., 2023c; Rajput et al., 2023). However, these methods only consider the item content information while the collaborative signals between users and items are overlooked. On the other hand, Si et al. (2023) construct GIDs using item embeddings of a pretrained SASRec (Kang and McAuley, 2018). Although the GID in this work contains the item-item sequential (collaborative) connections, the proposed recommendation framework in their research fails to model the item content information. To achieve satisfying recommendation results, we argue that the generative recommenders should be able to jointly model the item content information and the user-item collaborative signals in a unified generative framework.
@@
Secondly, existing generative recommendation methods cannot effectively align item content information and collaborative signals. Although Hua et al. (2023) proposed to construct the GID using the concatenation of a semantic string learnt from the item content and a sequence of tokens learnt from the item-item co-occurrence matrix, the naive concatenation without a learning process is not effective to perform the alignment, leading to sub-optimal performance. The rich content information can assist the recommendation models in obtaining more informative and comprehensive item representations, thus enhancing the modeling of user-item collaborative interactions, and vice visa. To this end, the alignment between collaborative signals and content information is fundamental for further enhancing generative recommenders. However, we argue that a successful alignment should be achieved through an explicit learning process, e.g.,the mapping between the content-based semantic space and the interaction-based collaborative space.
Proposed Methods. In this paper, we propose content-based colla-borative generation for recommender systems (ColaRec), a generative recommendation model which unifies both item content information and user-item collaborative interaction signals in a sequence-to-sequence generation framework. Taking a particular user as an example, the input sequence of ColaRec consists of unordered [^2] In this paper, we target on the general recommendation task other than sequential recommendation. To this end, we use unordered item tuples as the input. tuples with each tuple describing the content information of one interacted item of this user. Then an encoder-decoder framework is used to generate the GID of the target item.
@@
To address the first limitation, we propose two tailored designs to jointly model user-item collaborative signals and item content information. Firstly, we propose to construct GIDs using item representations obtained from a pretrained collaborative filtering model. In this paper, we use LightGCN (He et al., 2020) as the pretrained model. The LightGCN model is trained on the user-item interaction graph and thus the constructed GID can effectively encodes the user-item collaborative signals. Note that the LightGCN model can also be alternated with other models. Secondly, in ColaRec, the user is represented as the content aggregation of tuples with each tuple describing one her/his historically interacted item. This design also keeps inline with the nature of collaborative filtering. The aggregation of content-based tuples is fed to an encoder-decoder based language model to effectively capture the textual content information. In this paper, we use a fine-tuned T5 (Raffel et al., 2020) as the involved language model. We leave the investigation of larger language models as one of our future works.
@@
To address the second limitation, we propose an auxiliary item indexing task which targets on mapping the item side information into the GID of this item through the same encoder-decoder model. More precisely, the item side information contains both the textual content information and a set of users who have interacted with this item. To this end, the indexing task maps both the item content information and user-item interaction signals into the constructed GID, achieving better alignment between item content information and user-item collaborative signals. Besides, we further propose a contrastive loss to ensure that items with similar collaborative GIDs are also similar in the content-based semantic space. Such the contrastive loss also helps to perform alignment between content information and collaborative signals.
@@
To demonstrate the effectiveness of the proposed ColaRec, we conduct extensive experiments on three public accessible benchmark datasets. Experimental results show that the proposed ColaRec outperforms related state-of-the-art baselines.
Main Contributions. Our main contributions are as follows:
@@
  • We propose ColaRec, an end-to-end generative recommender which utilizes an encoder-decoder based language model to jointly model item content information and user-item collaborative signals for recommendation.
@@
  • We conduct extensive experiments on three datasets to demonstrate the effectiveness of the proposed ColaRec. Experimental results show superior recommendation performance of ColaRec.
@@
In this section, we review related prior literature on collaborative filtering and generative models.
@@ 2.1. Collaborative Filtering
Collaborative filtering (CF) is one of the most representative methods to build a recommendation agent. CF believes that a user can be represented as the aggregation of her/his interacted items, and vice visa. The keystone to conduct collaborative filtering is the user-item interaction matrix. Early approaches (Koren et al., 2009; Rendle et al., 2012) are based on matrix factorization (MF) to jointly model the latent space for users and items. Due to the expressiveness of deep neural networks, plenty of research (He et al., 2017; Mao et al., 2021; Zheng et al., 2017; Chen et al., 2018; Gao et al., 2020) has been conducted to enhance CF through deep learning. Besides, since the user-item interaction signals can be naturally encoded into an interaction graph, graph neural networks (GNN) also shed lights in the field of CF. Berg et al. (2017) proposed to use graph convolution for matrix completion. Wang et al. (2019) proposed the NGCF model to use GNN for CF. He et al. (2020) proposed the LightGCN model to simplify GNN for recommender systems, leading to a simpler and linear CF model. LightGCN serves as one of the most popular GNN-based CF approaches due to its effectiveness. Lin et al. (2022) proposed NCL to introduce contrastive learning into graph-based CF. Besides, content information has also been utilized to enhance the CF model. Rendle (2010) proposed the notable factorization machine (FM) to extend MF for categorical contextual features. He and McAuley (2016); Wei et al. (2019, 2020); Wu et al. (2017) proposed to enhance CF with visual or text features. Li et al. (2023b) proposed RecFormer to model long text sequences for recommendation.
Different with existing CF methods, in this paper we focus on the new paradigm of generative recommendation, where the recommended item is generated through the decoder of a sequence-to-sequence model. Generative recommendation provides new potential to utilize advances in the field of generative artificial intelligence to enhance the recommendation task.
2.2. Generative Models
@@
Generative models have become a hot research topic to generate new content, such as images and text, from the collected data. Variational Autoencoders (VAEs) (Kingma and Welling, 2014; Higgins et al., 2016) and Generative Adversarial Networks (GANs) (Goodfellow et al., 2014; Karras et al., 2019; Creswell et al., 2018), are two kinds of representative generative models. VAEs aim to learn a probabilistic mapping from the observed data to a latent space, and then leverage variational inference to approximate the posterior distribution of latent variables. GANs introduce a min-max gaming approach with a generator attempting to generate simulated samples, and the discriminator distinguishing between real and generated data. Recently, Transformers (Vaswani et al., 2017) have shown promise in language generation, leading to notable LLMs like GPTs (Generative Pre-trained Transformers). Besides, diffusion models (Ho et al., 2020; Yang et al., 2023; Cao et al., 2022) have also shown promising results in content generation.
Generative models for document retrieval. Generative retrieval has drawn increasing attention to conduct document retrieval with a sequence-to-sequence framework, where the input sequence is the query and the output sequence is the document identifier. The bloom of LLMs encourages generative retrieval to better understand semantic patterns stored in the retrieval process. Tay et al. (2022) firstly proposed a differentiable search index (DSI) for document retrieval, which assigns each document with a structured semantic meaningful identifier. DSI can be regarded as the early attempt to conduct generative retrieval. Plenty of work has been conducted following this research line. For example, strings including titles (Tang et al., 2023; Cao et al., 2021), n-grams (Bevilacqua et al., 2022; Chen et al., 2023) or URLs (Zhou et al., 2022) are explored to construct the document identifier. Wang et al. (2022a) proposed NCI which produces a hierarchical document identifier to get fine-grained semantic mapping, and a novel decoder to consider the prefixes of document identifiers. GenRet (Sun et al., 2023) learns to tokenize the document into a short discrete representation via a discrete auto-encoding approach, and utilizes progressive training and diverse clustering to assist the model training process.
Figure 3. Overview of ColaRec. ColaRec assigns each item with a GID obtained from a GNN-based CF model. ColaRec consists two tasks. User-item recommendation aims to map the user’s interacted items with textual content into the GID of the recommended item, i.e., $\mathcal{L}_{\text{rec}}$. Item-item indexing targets on the mapping from item side information into the item’s GID, i.e., $\mathcal{L}_{\text{index}}$. Besides, a ranking loss $\mathcal{L}_{\text{bpr}}$ and a contrastive loss $\mathcal{L}_{\text{c}}$ are also introduced. A tuned T5 is used to accomplish these tasks.
Generative models for recommendation. Generative models have also been utilized for recommender systems. Early works focus on utilizing VAEs (Liang et al., 2018; Shenbin et al., 2020; Cai and Cai, 2022) or GANs (He et al., 2018; Guo et al., 2022; Wang et al., 2022b) for recommendation. Recently, diffusion models have also been utilized to perform item recommendation (Wang et al., 2023b; Li et al., 2023a). Besides, LLM-based recommendation has also been proposed (Bao et al., 2023; Zhang et al., 2023; Petrov and Macdonald, 2023). These works explore the ability of LLMs for recommendation through parameter efficient tuning (Bao et al., 2023) or instruction tuning (Zhang et al., 2023). Meanwhile, several works attempt to convert various recommendation tasks into a unified natural language generation task, and train recommenders with multi-task optimization, like P5 (Geng et al., 2022) and M6-Rec (Cui et al., 2022).
Notably, TIGER (Rajput et al., 2023) is a representative generative recommendation method which uses an RQ-VAE (Zeghidour et al., 2022) to construct the GID and then uses encoder-decoder based transformers to generate sequential recommendation. Si et al. (2023) proposed to construct the GID from a pretrained SASRec (Kang and McAuley, 2018) model for sequential recommendation. Hua et al. (2023) further investigated the effect of item identifier construction for generative recommendation.
@@
Despite the arise of generative recommenders, existing methods still suffer from the ineffective infusion of collaborative signals and item content information. How to jointly model and align collaborative signals and content information in a sequence-to-sequence generative framework for recommender systems is still an open research challenge.

### v2

# Content-Based Collaborative Generation for Recommender Systems

@@ Abstract.
Generative models have emerged as a promising utility to enhance recommender systems. It is essential to model both item content and user-item collaborative interactions in a unified generative framework for better recommendation. Although some existing large language model (LLM)-based methods contribute to fusing content information and collaborative signals, they fundamentally rely on textual language generation, which is not fully aligned with the recommendation task. How to integrate content knowledge and collaborative interaction signals in a generative framework tailored for item recommendation is still an open research challenge.
In this paper, we propose content-based colla-borative generation for recommender systems, namely ColaRec. ColaRec is a sequence-to-sequence framework which is tailored for directly generating the recommended item identifier. Precisely, the input sequence comprises data pertaining to the user’s interacted items, and the output sequence represents the generative identifier (GID) for the suggested item. To model collaborative signals, the GIDs are constructed from a pretrained collaborative filtering model, and the user is represented as the content aggregation of interacted items. To this end, ColaRec captures both collaborative signals and content information in a unified framework. Then an item indexing task is proposed to conduct the alignment between the content-based semantic space and the interaction-based collaborative space. Besides, a contrastive loss is further introduced to ensure that items with similar collaborative GIDs have similar content representations. To verify the effectiveness of ColaRec, we conduct experiments on four benchmark datasets. Empirical results demonstrate the superior performance of ColaRec.
Recommender System, Generative Recommendation
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management; October 21–25, 2024; Boise, ID, USAbooktitle: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management (CIKM ’24), October 21–25, 2024, Boise, ID, USAdoi: 10.1145/3627673.3679692isbn: 979-8-4007-0436-9/24/10ccs: Information systems Recommender systems
@@
Recommender systems are widely deployed to discover user interests and provide personalized information services (Hu et al., 2018; Qi et al., 2020; Yuan et al., 2019). Recently, generative models, such as large language models (LLMs) (Zhao et al., 2023; Chang et al., 2023) and diffusion models (Yang et al., 2023; Cao et al., 2022), have gained prominence in advancing artificial intelligence. In such a context, plenty of research has emerged to utilize generative models for recommendation (Wang et al., 2023b; Rajput et al., 2023). To achieve satisfying results, the recommender should be able to incorporate both user-item collaborative signals and item content information. Collaborative signals refer to the knowledge contained in the user-item interactions while item content information refers to the textual description of items, as shown in Figure 1.
LLM-based recommendation is a straightforward solution to fuse content information with collaborative signals. In general, these approaches reorganize the historical user-item interactions into text-based natural language, aka the instruction-tuning dataset. Then, the recommendation task is reformulated as language generation under designed prompts  (Fan et al., 2023; Li et al., 2024). Although this kind of approach utilizes the powerful text generation capability of LLMs, there exists the inherent misalignment between the two tasks of language generation and item recommendation. For example, there is usually a non-trivial grounding stage to map the generated language to a concrete item (Lin et al., 2023; Bao et al., 2023). LLMs also suffer poor ranking performance to generate target itemIDs from a large candidate pool (Zhang et al., 2023b; Liao et al., 2023). The task discrepancy limits the practical usage of LLM-based methods.
Inspired by generative retrieval (Tay et al., 2022), some works (Rajput et al., 2023; Si et al., 2023) assign each item with a unique sequence of tokens as the item’s generative identifier, aka GID, then a generative sequence-to-sequence model which is tailored for directly generating item recommendation is trained without the need of explicit language modeling. The input sequence comprises historical user-item interactions while the output sequence refers to the GID of the recommended item. In this paper, we refer such the above paradigm as Generative Recommendation [^1] Wang et al. (2023a) proposed another paradigm to directly generate new content, e.g., images, for recommendation. However, their methods are tailored for the generation of virtual content and cannot be used for concrete items. In this paper, we target on the recommendation of concrete items.. Compared with conventional itemIDs with an assigned single random token, the sequential tokens of a GID contain more explicit representation information. As shown in Figure 2, we see correlated GIDs denote correlated items, and thus help the recommender to conduct more effective generation. Generative recommendation provides an end-to-end paradigm to design generative models tailored for the recommendation task. Plenty of studies have been conducted along this research line  (Rajput et al., 2023; Si et al., 2023; Petrov and Macdonald, 2023; Tan et al., 2024; Zheng et al., 2023).
However, existing generative recommendation methods still fail to effectively model collaborative signals and content information in a unified framework. For example, several methods utilize either the item titles or hierarchical content embeddings of items to construct GIDs (Li et al., 2023c; Rajput et al., 2023). These methods only consider the item content information while the collaborative signals between users and items are overlooked. On the contrary, Si et al. (2023) constructed GIDs using item embeddings of a pretrained SASRec (Kang and McAuley, 2018). Although the GID in this work contains the item-item sequential (collaborative) connections, the proposed recommendation framework in their research fails to model the item content information.
@@
Besides, existing generative recommendation methods cannot effectively align item content information and collaborative signals. Although Hua et al. (2023) proposed a combined approach that concatenates content-based semantic strings with collaborative IDs derived from the item-item co-occurrence matrix. However, this naive concatenation, lacking a proper learning process, fails to achieve effective alignment, resulting in sub-optimal performance. The alignment should be achieved through an explicit learning process, e.g., the mapping between the content-based semantic space and the interaction-based collaborative space.
@@
In this paper, we propose content-based colla-borative generation for recommender systems (ColaRec), a model that unifies both item content information and user-item collaborative signals in a sequence-to-sequence generation framework tailored for the recommendation task. Taking a particular user as an example, the input sequence of ColaRec consists of unordered [^2] In this paper, we focus on the general recommendation task other than sequential recommendation. To this end, we use unordered item tuples as the input. tuples with each tuple describing the content information of one interacted item of this user. Then an encoder-decoder Transformer (Vaswani et al., 2017) is used to generate the GID of the target item. More precisely, we propose two tailored designs to jointly model user-item collaborative signals and item content information. Firstly, we propose to construct GIDs using item representations obtained from a pretrained collaborative filtering model. In this paper, we use LightGCN (He et al., 2020) as the pretrained model. The LightGCN model is trained on the user-item interaction graph and thus the constructed GID can effectively encode the user-item collaborative signals. Note that the LightGCN model can also be alternated with other models. Secondly, in ColaRec, the user is represented as the content aggregation of tuples with each tuple describing one her/his historically interacted item. This design also keeps inline with the nature of collaborative filtering. The aggregation of content-based tuples is fed to the generative Transformer model to effectively capture the item content information.
@@
To conduct the alignment between content information and collaborative signals, we propose an auxiliary item indexing task which targets on mapping the item side information into the GID of this item through the same encoder-decoder model. Specifically, the item side information contains both the textual content information and a set of users who have interacted with this item. To this end, the indexing task maps both the item content information and user-item interaction signals into the constructed GID, achieving better alignment. Besides, we further propose a contrastive loss to ensure that items with similar collaborative GIDs are also similar in the content-based semantic space.
@@
To demonstrate the effectiveness of the proposed ColaRec, we conduct extensive experiments on four public datasets. Experimental results show that the proposed ColaRec outperforms related state-of-the-art baselines [^3] The code of this work is available at https://github.com/Junewang0614/ColaRec.
Our main contributions are as follows:
@@
  • We propose ColaRec, a generative recommendation framework which utilizes an encoder-decoder model to jointly capture content information and collaborative signals for recommendation.
@@
  • We conduct extensive experiments on four datasets to demonstrate the effectiveness of the proposed ColaRec. Experimental results show superior recommendation performance of ColaRec.
@@
In this section, we review related literature on collaborative filtering and generative models for recommendation.
@@ 2.1. Collaborative Filtering
Collaborative filtering (CF) is one of the most representative methods to build a recommendation agent. CF believes that a user can be represented as the aggregation of her/his interacted items, and vice versa. The keystone to conduct collaborative filtering is the user-item interaction matrix. Early approaches (Koren et al., 2009; Rendle et al., 2012) are based on matrix factorization (MF) to jointly model the latent space for users and items. Due to the expressiveness of deep neural networks, plenty of research (He et al., 2017; Mao et al., 2021; Zheng et al., 2017; Chen et al., 2018; Gao et al., 2020) has been conducted to enhance CF through deep learning. Besides, since the user-item interaction signals can be naturally encoded into an interaction graph, graph neural networks (GNN) also shed lights in the field of CF. Berg et al. (2017) proposed to use graph convolution for matrix completion. Wang et al. (2019) proposed the NGCF model to use GNN for CF. He et al. (2020) proposed the LightGCN model to simplify GNN for recommender systems, leading to a simpler and linear CF model. LightGCN serves as one of the most popular GNN-based CF approaches due to its effectiveness. Lin et al. (2022) proposed NCL to introduce contrastive learning into graph-based CF. Besides, content information has also been utilized to enhance the CF model. Rendle (2010) proposed the notable factorization machine (FM) to extend MF for categorical contextual features. He and McAuley (2016); Wei et al. (2019, 2020); Wu et al. (2017) proposed to enhance CF with visual or text features. Li et al. (2023b) proposed RecFormer to model long text sequences for recommendation.
Different from existing CF methods, in this paper we focus on the new paradigm of generative recommendation, where the item is represented as a sequence of tokens, i.e., GID, and the recommendation is provided in a generative fashion.
2.2. Generative Models for Recommendation
@@
Generative models have become a hot research topic to generate new content, such as images and text. Variational Autoencoders (VAEs) (Kingma and Welling, 2014; Higgins et al., 2016) and Generative Adversarial Networks (GANs) (Goodfellow et al., 2014; Karras et al., 2019; Creswell et al., 2018), are two kinds of representative generative models. Besides, diffusion models (Ho et al., 2020; Yang et al., 2023; Cao et al., 2022) have also shown promising results in content generation. These generative models have also been utilized for recommender systems  (Liang et al., 2018; Shenbin et al., 2020; Cai and Cai, 2022; He et al., 2018; Guo et al., 2022; Wang et al., 2022b, 2023b; Li et al., 2023a).
Recently, Transformers (Vaswani et al., 2017) have shown promise in language generation, leading to notable LLMs like GPTs (Generative Pre-trained Transformers). LLM-based recommendation provides a straightforward solution to fuse user-item interaction signals with item textual content  (Liao et al., 2023; Zheng et al., 2023; Zhang et al., 2023a; Luo et al., 2024; Zhu et al., 2023). Typically, these approaches need to construct the recommendation training data with natural language, i.e., the instruction-tuning dataset, then the recommendation task is transformed into text generation under designed language prompts  (Zhang et al., 2023b; Lin et al., 2023; Fan et al., 2023; Li et al., 2024). For instance, Luo et al. (2024) proposed to inject collaborative signals into LLM-based recommendation through prompt augmentation. (Zhang et al., 2023a) and (Liao et al., 2023) proposed to fuse item embeddings into prompt embeddings for CTR prediction and recommendation. Lin et al. (2023) utilized LLMs to generate multiple aspects of the recommended items including titles and attributes. LC-Rec (Zheng et al., 2023) proposed to perform recommendation through conducting various language generation tasks. Despite the strong generation capabilities of LLMs, there is an inherent misalignment between language generation and item recommendation, which limits the practical usage of LLM-based methods. For example, it is usually necessary to go through a non-trivial grounding stage to map the generated language into specific items (Bao et al., 2023; Lin et al., 2023; Liao et al., 2023). Besides, it is also challenging for LLMs to directly generate recommended items from a large candidate pool (Liao et al., 2023; Zhang et al., 2023b; Luo et al., 2024).
Inspired by generative retrieval, generative recommendation is a new paradigm to design generative models which is tailored for the recommendation task. Generative retrieval, which relies on the parametric memory of generative models to directly generate relevant document identifiers, has drawn increasing attention. Tay et al. (2022) firstly proposed a differentiable search index (DSI) for generative retrieval. Plenty of methods have been conducted following this research line. Representative works include SEAL (Bevilacqua et al., 2022), NCI (Wang et al., 2022a), and GenRet (Sun et al., 2023). Similar to generative retrieval, generative recommendation constructs a sequence of tokens, aka generative identifier (GID), for each item, then a generative sequence-to-sequence model which is tailored for directly generating item recommendation is trained without the need of explicit language modeling. The input comprises historical user-item interactions and the output refers to the GID of the recommended item. Following this paradigm, TIGER (Rajput et al., 2023) is a representative generative recommendation method which uses an RQ-VAE (Zeghidour et al., 2022) to construct the GID and then uses a encoder-decoder based transformer to generate sequential recommendation. Si et al. (2023) proposed to construct the GID from a pretrained SASRec (Kang and McAuley, 2018) model for sequential recommendation. Hua et al. (2023) further investigated the effect of item identifier construction. Tan et al. (2024) utilized human language tokens to represent each item with a unique textual identifier.
@@
Despite the rise of generative recommendation, existing generative recommendation methods still suffer from the ineffective infusion of collaborative signals and item content information. How to jointly model and align collaborative signals and content information in an end-to-end generative framework tailored for item recommendation is still an open research challenge.
Figure 3. Overview of ColaRec. ColaRec assigns each item with a GID obtained from a GNN-based CF model. ColaRec consists two tasks. User-Item Recommendation aims to map the user’s interacted items with textual content into the GID of the recommended item, i.e., $\mathcal{L}_{\text{rec}}$. Item-Item Indexing targets on the mapping from item side information into the item’s GID, i.e., $\mathcal{L}_{\text{index}}$. Besides, a ranking loss $\mathcal{L}_{\text{bpr}}$ and a contrastive loss $\mathcal{L}_{\text{c}}$ are also introduced.